### v1

# An Embedding is Worth a Thousand Noisy Labels

Abstract
The performance of deep neural networks scales with dataset size and label quality, rendering the efficient mitigation of low-quality data annotations crucial for building robust and cost-effective systems. Existing strategies to address label noise exhibit severe limitations due to computational complexity and application dependency. In this work, we propose WANN, a Weighted Adaptive Nearest Neighbor approach that builds on self-supervised feature representations obtained from foundation models. To guide the weighted voting scheme, we introduce a reliability score $\eta$, which measures the likelihood of a data label being correct. WANN outperforms reference methods, including a linear layer trained with robust loss functions, on diverse datasets of varying size and under various noise types and severities. WANN also exhibits superior generalization on imbalanced data compared to both Adaptive-NNs (ANN) and fixed $k$-NNs. Furthermore, the proposed weighting scheme enhances supervised dimensionality reduction under noisy labels. This yields a significant boost in classification performance with 10×10$\times$10 × and 100×100$\times$100 × smaller image embeddings, minimizing latency and storage requirements. Our approach, emphasizing efficiency and explainability, emerges as a simple, robust solution to overcome the inherent limitations of deep neural network training. The code is available at github.com/francescodisalvo05/wann-noisy-labels.
keywords:
Label noise, Efficiency, k-NN, Adaptive-NN, Weighted-NN, Robust loss functions
Figure 1: Illustration of the proposed Weighted Adaptive Nearest Neighbor (WANN) algorithm. Initially, a reliability score ($\eta$) is computed for each training observation, representing the inverse of the minimum number of samples needed for a correct prediction. During inference, the adaptive neighborhood size ($k_{T}$) of each test observation ($x_{T}$) is determined based on the reliability score of its closest training point ($k_{T}=\frac{1}{\eta}=3$). A weighted majority vote determines the final label, reducing the impact of noisy labels.
1 Introduction
The remarkable results achieved by deep neural networks in a multitude of domains are possible due to ever-increasing computational power. This progress has enabled the development of deeper architectures with stronger learning capabilities. However, these high-parametric architectures are often characterized as data-hungry, because they require a large amount of data to generalize effectively. Collecting and annotating such extensive datasets can be both expensive and time-consuming, potentially introducing machine and human errors. In fact, the proportion of corrupted labels in real-world datasets ranges from 8-38.5% [42]. While unsupervised and semi-supervised methods have garnered significant attention within the research community lately [7, 44, 1], supervised methods are still widely employed due to their generally higher performance. In this context, incorrect labels pose significant challenges, including a tendency to memorize label noise, negatively affecting generalization capabilities. This issue, as demonstrated by [56], is hindering the adoption of AI systems in safety-critical domains, such as healthcare.
Consequently, there is a growing interest in enhancing the robustness of deep models against noisy labels, resulting in five different lines of research [42]: robust architectures, robust regularization, robust loss function, loss adjustment and sample selection. Nevertheless, as also highlighted by [58], all these approaches involve a learning process and, by definition, have limitations in generalizing to diverse datasets or varying noise rates.
We address those limitations and the challenges posed due to noisy labels by exploiting feature representations obtained from large pre-trained models, commonly referred to as foundation models. Many of these foundation models are now publicly available and designed for a wide range of applications, including image classification, semantic segmentation tasks, and beyond. While their initial focus was primarily on text [12, 6] and natural images [11, 9, 34, 38], there is currently a focus on the development of open-source foundation models tailored to specific domains, such as healthcare [43, 29, 57, 52], making it possible to translate this paradigm to various applications with little effort. Specifically, large image-based foundation models are usually trained in a self-supervised fashion, e.g., by means of contrastive learning [7], representing similar objects closely within the embedding space, and semantically distinct objects further away. Thus, the position of an image in the embedding space can be just as informative as the label, if not more so. Although embedding-space approaches are not new in the literature to process noisy labels, they typically focus on noise detection, through online [5] or offline [58] methods. However, thanks to the representational power of modern embedding spaces [38, 34], we believe that it is also possible to provide robust predictions by building on simple $k$-NN methods. This category of approaches not only offers computational efficiency but also enhances explainability. Moreover, the lower reliance on hyperparameters further enhances generalizability. Thus, owing to its simplicity and efficiency, it is able to mitigate some of the limitations associated with the training of deep networks.
In this work, we propose a Weighted Adaptive Nearest Neighbor (WANN) approach, illustrated in Figure 1, which operates on the image embeddings extracted from a pre-trained foundation model. The introduced weighted adaptive voting scheme addresses the challenges posed by large-scale, limited, and imbalanced noisy datasets. Our contributions are:
  • Formulation of a reliability score ($\eta$), measuring the likelihood of a label being correct. This scoring guides the formulation of WANN, a method that determines a neighborhood of the test sample to weight an adaptively determined number of training samples in the embedding space.
  • Extensive quantitative experiments, confirming that WANN has overall greater robustness compared to reference methods (ANN, fixed $k$-NN, robust loss functions) across diverse datasets and noise levels, including limited and severely imbalanced noisy data scenarios.
  • Formulation of the Filtered LDA (FLDA) approach that relies on our sample-specific reliability scores for higher quality projections. FLDA improves the robustness for dimensionality reduction by filtering out detected noisy samples and improves the classification performances with 10×10$\times$10 × and 100×100$\times$100 × smaller image embeddings.
Consequently, through lightweight embeddings and a simple yet efficient classification algorithm, we demonstrate that working efficiently within the embedding space might constitute a potential paradigm shift, increasing efficiency and robustness while addressing several critical limitations associated with training deep neural networks.
((a)) Wrong test labels
((b)) Ambiguous test labels
Figure 2: Example of test images and their relative top-$3$ closest training samples extracted from the STL-10 dataset [8]. Figure 2(a) shows two test images having wrong labels, while Figure 2(b) shows ambiguous labels, including multiple known objects within a single image. For instance, a bird with a ship in the background, or a dog fighting with a cat.
2 Related works
Noise robust learning   Robust learning from noisy labels has received considerable attention in recent years, and multiple deep-learning-based methods are currently available [42]. Effective and popular approaches involve multi-network learning, where noisy samples during training are discarded. The sample selection might be guided by a mentor network [23] or a peer network [19, 53, 46]. However, due to their computational complexity, these models often lack generalizability across datasets or noise settings. Some of the lightweight techniques designed to handle noisy labels fall under the category of robust loss functions. While the Cross Entropy (CE) suffers from overfitting to wrong labels [56], the Mean Absolute Error (MAE) has been theoretically guaranteed to be noise-tolerant [15]. However, it suffers from severe underfitting in challenging domains. To address this issue and enhance its generalization, the Generalized Cross Entropy (GCE) [60] was introduced as a generalization of both MAE and CE. While [47] propose the Symmetric Cross Entropy (SCE), [59] overcome the symmetric condition through Asymmetric Loss Functions. Among state-of-the-art methods, we recall Early-Learning Regularization [28] which tries to prevent memorization of the false labels via regularization, and Active Passive Losses (APL) [31] which combines two robust losses to balance between overfitting and underfitting. Recently, Active Negative Losses (ANL) [54] replaced the previous passive loss (in APL) with a Normalized Negative Loss Function (NNLF). Nevertheless, despite their relatively lower computational complexity compared to multi-network methods, they still inherit challenges related to training neural networks. Some of these challenges include the demand for large datasets, lack of explainability, computational complexity, hyperparameter dependence, and overfitting towards wrong labels.
$k$-NN for label noise   Besides native deep learning-based approaches, certain traditional machine learning methods are robust by design. Notably, $k$-Nearest-Neighbors-based methods have recently gained attention for their utility in filtering out noisy training observations [37, 5, 26]. Although the proposed online $k$-NN approaches can enhance the robustness of DNNs trained with label noise, it does not address the fundamental issues associated with DNNs, such as the lack of explainability, data-efficiency, and generalizability. A closely related work [58] already emphasized the potential of a training-free cleaning strategy based on $k$-NN, relying on CLIP [38] embeddings. In alignment with this research direction, we aim to accentuate the intrinsic robustness and explainability benefits (cf. Figure 2) of $k$-NN in classification tasks. This is further enhanced by the adoption of an adaptive neighborhood for each test observation, defined through a reliability scoring.
3 Method
At its core, the proposed method relies on an adaptive $k$-NN search over the image embeddings extracted from a pre-trained foundation model, using the Euclidean distance as reference distance metric. In particular, we employ the DINOv2 [34] Large backbone, with $14{\times}14$ patches, yielding image embeddings of size $1024$. DINOv2 was trained in a self-supervised fashion on $142$M images, and the choice of the backbone is further motivated in Section 4.1. To enhance computational efficiency, we pre-generated a database of embeddings for all of the evaluated datasets, following the approach adopted by [33].
Label reliability score   Although $k$-NN is robust against noisy labels by design, its efficacy may diminish under severe noise conditions. To address this challenge, the introduction of a weighting scheme becomes helpful, aiming to mitigate the impact of noisy labels. Intuitively, a proper weighting scheme should assign higher weights to samples that are more likely to have a correct label, and lower weights to those potentially mislabeled. In this context, we introduce the concept of label reliability, building upon the work of [40]. Extending their approach, which determined an optimal $k$ value for each training observation, we determine the reliability (i.e., quality) of the associated label, guiding a subsequent weighting scheme. For a given training observation $x_{i}$ and its associated label $y_{i}$, the reliability $\eta_{i}$ is defined as the inverse of the minimum number of training samples (neighborhood size) required for a correct $k$-NN classification. Dropping the index $i$ for readability, we express $\eta$ as the function $\mathcal{H}(x,y;\mathcal{X})$, where $\mathcal{X}$ represents the training dataset:
$\mathcal{H}(x,y;\mathcal{X})=\frac{1}{\underset{{k\in[k_{\text{min}},k_{\text{max}}]}}{\arg\,\min}{\,}{\texttt{kNN}_{\mathcal{X}}(x,k)=y}}$ (1)
Here $k_{\text{min}}$ and $k_{\text{max}}$ define the search interval for $k\in\mathbb{N}$. It was previously shown by [58] that for a fixed $k$-NN approach a suitable value for $k$ is $10$ to detect corrupted labels. Consequently, we set $k_{\text{min}}=11$ and $k_{\text{max}}=51$. This allows to establish a substantial search space covering a wide range of noise levels while ensuring a baseline ability of detecting corrupt labels ($k>10$). During parameter search we increase $k$ by $2$ at each incremental step, ensuring an odd number of samples in the neighborhood while reducing the computational complexity.
The correlation between the value of $k$ and the reliability of the assigned label stems from the main property of the embedding space. The proximity of similar objects suggests that a small $k$ should be sufficient for accurate classification. In such cases, the reliability of the data point is higher ($k\downarrow\;\eta\uparrow$). Conversely, if a data point requires a large $k$ value for correct classification, approaching or reaching $k_{\text{max}}$, it may signal potential mislabeling ($k\uparrow\;\eta\downarrow$). The pseudocode is presented in Algorithm 1, and it requires three parameters: $k_{\text{min}},k_{\text{max}}$, and $D_{\text{train}}$. The latter, $D_{\text{train}}$, represents a data structure containing a triple $(\texttt{key},x,y)$ for each train observation. Here, key serves as a unique identifier.
Algorithm 1 reliability($k_{\text{min}},k_{\text{max}},D_{\text{train}}$)
[…]
  $P\leftarrow\texttt{PairwiseDist}(D_{\text{train}},D_{\text{train}})$matrix}
  for $(\texttt{key},x,y)$ in $D_{\text{train}}$ do
     for $k=k_{\text{min}},k_{\text{min}}+2,\ldots,k_{\text{max}}$ do
        $\hat{y}\leftarrow\texttt{kNN}_{P}(x,k)$$k$ points excluding $x$}
        if $\hat{y}==y$ then
           $R[\texttt{key}]\leftarrow\frac{1}{k}$
           break
     if not key in R.keys() then
        $R[\texttt{key}]\leftarrow\frac{1}{k_{\text{max}}}$
  return  $R$
Weighted Adaptive Nearest Neighbor (WANN)   Adaptive Nearest Neighbors algorithms were commonly employed for tabular data with limited number of observations and classes [16, 45, 48]. Following the intuition of [40], we use $\eta$ to adaptively determine the neighborhood size for any test observation. In fact, if a test sample is close to a potentially noisy label ($\eta\downarrow$) a larger neighborhood is preferable to reduce variance. Conversely, if a test observation is close to a possibly clean label ($\eta\uparrow$), a small neighborhood increases specificity. Thus, given a test observation $x_{T}$ with nearest training sample $(x_{n},y_{n},\eta_{n})$ (nearest in the embedding space), its test neighborhood size $k_{T}$ will be adaptively defined by means of the reliability score of $x_{n}$ (cf. Equation 2).
$k_{T}=\frac{1}{\eta_{n}}=k_{n}$ (2)
With this notation we formulate an Adaptive Nearest Neighbor (ANN) method as baseline, similar to [40]:
$\texttt{ANN}(x_{T})=\underset{c\in C}{\arg\,\max}\frac{1}{k_{T}}\sum_{i\in N_{T}}\mathds{1}(y_{i}=c)$ (3)
Here, $C$ denotes the set of classes, $N_{T}$ refers to the adaptive neighborhood determined by the $k_{T}$ closest observations, and $\mathds{1}(\cdot)$ is the indicator function. Finally, having defined the reliability score as ${\eta=\mathcal{H}(x,y;\mathcal{X})}$, we enhance ANN by introducing a weighting scheme. Thus, instead of merely taking the most frequent class in the neighborhood, we now associate a weight (i.e., $\eta$) to each training observation. Consequently, training samples with higher reliability $\eta$ will have a greater impact on the final prediction, while those with lower $\eta$ (possibly wrong labels) will contribute less. As such, our Weighted Adaptive Nearest Neighbor approach is formulated as:
$\texttt{WANN}(x_{T})=\underset{c\in C}{\arg\,\max}\sum_{i\in N_{T}}\eta_{i}\,\mathds{1}(y_{i}=c)$ (4)
[…]
To address this limitation, we employ WANN’s reliability score $\eta$ to filter out potentially wrong labels (i.e., $\eta=1/{k_{max}}$) before performing LDA. We term this approach Filtered LDA (FLDA). Once the projection axes have been obtained through the clean dataset, we subsequently project both training and test datasets onto those axes.
4 Experiments and results
Initially, in Section 4.1, we justify the choice of the DINOv2 Large backbone [34]. Next, in Sections 4.2 – 4.5, we evaluate WANN against robust loss functions on noisy real-world, limited, and medical data. In Section 4.6, we assess the robustness of the weighted adaptive neighborhood in the context of heavily imbalanced datasets. We further demonstrate the effectiveness of the proposed Filtered LDA (FLDA) approach in Section 4.7. Finally, in Section 4.8, we qualitatively explore the explainability benefits of our approach on noisy datasets. It is important to note that none of the datasets used in our quantitative experiments were part of DINOv2’s training data, as detailed in Table 15 of [34].
4.1 Backbone
In order to achieve satisfactory $k$-NN performance, a high-quality feature representation is essential to ensure close proximity of similar objects. To evaluate this, we initially compare the performance of ResNet50111resnet50.tv_in1k and ResNet101222resnet101.tv_in1k [20], both pre-trained on ImageNet-1k. In the realm of Vision Transformers, we explore self-supervised pre-training with MAE [18] (Base333vit_base_patch16.mae and Large444vit_large_patch16.mae), CLIP pre-training [38] (Base555vit_base_patch16_clip_224.openai and Large666vit_large_patch14_clip_336.openai), and lastly DINOv2 pre-training [34] (Base777vit_base_patch14_dinov2.lvd142m and Large888vit_large_patch14_dinov2.lvd142m). All pre-trained models are obtained from HuggingFace (timm). We report WANN’s classification accuracy on CIFAR-10 and CIFAR-100 [25].
Results   Table 1 displays the results, clearly confirming DINOv2 as backbone choice. Results further reveal a notable decline in performance for all backbones as the number of classes increases. This trend is expected due to increasing overlaps of class-specific feature distributions within the embedding space. Additionally, obtaining ground truth labels for a larger set of visually more ambiguous classes poses more challenges associated with label quality. Notably, comparing MAE ViTs with ResNets, both pre-trained on ImageNet, it is possible to observe a substantially lower performance of MAE ViTs. Despite the higher number of parameters, the self-supervised pre-training paradigm appears to be not beneficial in this context. Furthermore, the discrepancy between CLIP and DINOv2 can be attributed to the divergent pre-training procedures: CLIP is trained on pairs of image-text, whereas DINOv2 exclusively relies on image data. Beyond computational efficiency, the DINOv2 framework allows to disregard potential ambiguities introduced by textual descriptions. For this reason, it also required considerably less data, as DINOv2 was trained on $142$ million images while CLIP on $400$ millions of image-text pairs. Thus, in scenarios where text embeddings are not explicitly needed, as in our case, opting for an image-centric training paradigm seems beneficial.
Backbone #Par #Dim C-10 C-100
ResNet50 26M 2048 84.09 60.31
ResNet101 45M 2048 87.32 64.59
MAE ViT-B/16 86M 768 72.04 45.18
MAE ViT-L/16 303M 1024 82.99 56.18
CLIP ViT-B/16 86M 768 92.24 68.96
CLIP ViT-L/14 304M 1024 95.65 75.69
DINOv2 ViT-B/14 87M 768 98.70 89.36
DINOv2 ViT-L/14 304M 1024 99.36 91.68
Table 1: WANN’s classification performance (accuracy) over eight different feature spaces, on CIFAR-10 and CIFAR-100, which contains 10 and 100 classes, respectively. Note that “#Par” denotes the number of parameters and “#Dim” denotes the feature dimension.
CIFAR-10N CIFAR-100N
Noisy split Clean Aggr. R1 R2 R3 Worst Clean Noisy
NR ($\approx$) - 9.01% 17.23% 18.12% 17.64% 40.21% - 40.20%
CE 99.40$\pm$0.03 98.46$\pm$0.07 98.29$\pm$0.11 98.33$\pm$0.11 98.45$\pm$0.05 93.20$\pm$0.32 93.38$\pm$0.13 83.10$\pm$0.29
FL 99.39$\pm$0.02 98.36$\pm$0.08 98.12$\pm$0.12 98.21$\pm$0.07 98.20$\pm$0.13 92.85$\pm$0.18 93.30$\pm$0.10 82.03$\pm$0.32
MAE 99.42$\pm$0.02 99.21$\pm$0.02 99.22$\pm$0.06 99.17$\pm$0.03 99.20$\pm$0.06 95.79$\pm$0.41 93.40$\pm$0.10 86.08$\pm$0.25
GCE 99.39$\pm$0.05 99.11$\pm$0.01 99.06$\pm$0.03 98.96$\pm$0.08 99.03$\pm$0.05 95.19$\pm$0.43 93.45$\pm$0.07 85.74$\pm$0.32
SCE 99.42$\pm$0.02 99.14$\pm$0.03 99.11$\pm$0.06 99.07$\pm$0.08 99.12$\pm$0.05 95.52$\pm$0.53 93.40$\pm$0.14 83.15$\pm$0.33
ELR 99.44$\pm$0.02 99.13$\pm$0.10 99.09$\pm$0.08 99.14$\pm$0.07 99.14$\pm$0.04 97.72$\pm$0.37 93.34$\pm$0.04 86.10$\pm$0.22
NCE-MAE 99.42$\pm$0.01 99.17$\pm$0.02 99.15$\pm$0.04 99.17$\pm$0.02 99.15$\pm$0.05 95.35$\pm$0.35 93.46$\pm$0.08 85.44$\pm$0.40
NCE-RCE 99.42$\pm$0.02 99.21$\pm$0.03 99.21$\pm$0.07 99.15$\pm$0.04 99.19$\pm$0.08 95.58$\pm$0.35 93.47$\pm$0.08 86.08$\pm$0.08
NFL-RCE 99.42$\pm$0.02 99.21$\pm$0.03 99.21$\pm$0.07 99.15$\pm$0.04 99.19$\pm$0.08 95.58$\pm$0.35 93.46$\pm$0.07 86.05$\pm$0.11
NCE-AGCE 99.42$\pm$0.02 99.21$\pm$0.03 99.21$\pm$0.05 99.18$\pm$0.03 99.18$\pm$0.06 95.93$\pm$0.43 93.40$\pm$0.07 86.25$\pm$0.12
NCE-AUL 99.42$\pm$0.02 99.20$\pm$0.03 99.21$\pm$0.06 99.18$\pm$0.04 99.18$\pm$0.05 95.77$\pm$0.48 93.41$\pm$0.09 85.81$\pm$0.27
NCE-AEL 99.41$\pm$0.02 99.14$\pm$0.01 99.08$\pm$0.06 99.04$\pm$0.09 99.11$\pm$0.04 95.33$\pm$0.63 93.43$\pm$0.09 84.79$\pm$0.38
ANL-CE 99.39$\pm$0.03 99.24$\pm$0.04 99.19$\pm$0.07 99.20$\pm$0.04 99.26$\pm$0.03 98.18$\pm$0.22 91.08$\pm$1.05 84.34$\pm$0.51
ANL-FL 99.39$\pm$0.04 99.25$\pm$0.04 99.21$\pm$0.04 99.19$\pm$0.03 99.24$\pm$0.03 98.23$\pm$0.24 90.77$\pm$0.20 84.19$\pm$0.44
ANN 99.37 99.19 99.10 99.06 98.97 95.31 91.99 84.91
WANN 99.36 99.32 99.27 99.19 99.24 97.21 91.68 85.44
Table 2: Accuracy ($\uparrow$) on two clean and noisy datasets [49]. The two best results are highlighted in bold. Note that NR stands for Noise Rate. CIFAR-10N includes three annotations per image, and Noisy split denotes various aggregation strategies. Following the proposed naming convention, Aggr refers to a majority voting, R-$i$ ($i\in\{1,2,3\}$) denotes the $i$-th submitted label for each image, and Worst denotes the selection of only wrong labels, if applicable. Conversely, CIFAR-100N presents just one annotation per image, thereby offering one single noisy split.
4.2 Real-world noisy labels
We compared ANN and WANN with lightweight techniques falling into the category of robust loss functions. As commonly employed for assessing the feature space quality [34], we trained a single linear layer over the image embeddings. The embeddings were normalized using training set statistics, excluding 15% for validation in linear training. The linear probing utilizes the Adam optimizer with a learning rate of $1{\times}10^{-4}$ for $100$ epochs, with early stopping after 15 epochs. These settings were consistently applied to all experiments in this manuscript. Following the publicly available losses and hyperparameters proposed by [54], we compared: Cross Entropy (CE) as a baseline, Focal Loss (FL), Mean Absolute Error (MAE), Generalized Cross Entropy (GCE), Symmetric Cross Entropy (SCE), Active Passive Losses (NCE+MAE, NCE+RCE, NFL+RCE), Asymmetric Losses (NCE+AGCE, NCE+AUL, NCE+AEL), and Active Negative Losses (ANL+CE, ANL+FL). Additionally, we include Early Learning Regularization (ELR), using the suggested hyperparameters. Since the embeddings are pre-generated, no data augmentation was applied. Both ANN and WANN use $(k_{\text{min}},k_{\text{max}})=(11,51)$. We measure the accuracy on two publicly available datasets, namely CIFAR-10N and CIFAR-100N [49], which are “noisy versions” of the established CIFAR datasets. CIFAR-10N contains three human annotations per image, proposing different noisy training splits with varying levels of noise rates (NR). Conversely, CIFAR-100N presents solely one annotation per image, thereby offering one single noisy split. Due to the potential sensitivity of linear training to the choice of the random seed, we show the accuracy as the mean and standard deviation computed over five runs.
Results   Table 2 shows the results of the designed experiment for real-world noisy labels. The introduced weighting scheme in WANN substantially enhances robustness compared to ANN, particularly in high-noise conditions. Additionally, WANN exhibits greater advantages compared to robust loss functions in scenarios with limited noise ratios (${<}20\%$), showing a notable lower performance drop than robust loss functions on CIFAR-10N Aggr and R1, respectively. Active Negative Losses (ANL-CE and ANL-FE) rank among the top robust loss functions, approaching WANN’s performance. Furthermore, while NCE-AGCE exhibits a higher accuracy on CIFAR-100N compared to WANN, it notably presents a higher drop in performance with respect to the clean performance.
CIFAR-10 (#$\#50$) CIFAR-10 (#$\#100$)
Pattern Clean Symmetric Asymmetric Instance Clean Symmetric Asymmetric Instance
NR - 60% 30% 40% - 60% 30% 40%
CE 93.09$\pm$1.40 53.40$\pm$10.31 75.75$\pm$6.62 66.34$\pm$21.03 95.78$\pm$1.14 74.10$\pm$3.42 85.22$\pm$1.66 87.83$\pm$2.23
FL 91.37$\pm$2.53 45.12$\pm$14.39 69.33$\pm$8.20 63.22$\pm$18.79 95.95$\pm$1.35 73.35$\pm$4.14 82.80$\pm$0.94 86.79$\pm$1.42
MAE 93.03$\pm$3.24 61.84$\pm$10.20 78.58$\pm$4.42 77.22$\pm$19.97 96.94$\pm$0.78 90.71$\pm$2.83 93.58$\pm$2.11 93.23$\pm$1.83
GCE 92.59$\pm$3.25 57.24$\pm$19.20 78.30$\pm$2.58 83.63$\pm$4.79 96.52$\pm$0.43 84.00$\pm$3.85 93.68$\pm$1.50 91.47$\pm$4.04
SCE 93.02$\pm$3.08 61.29$\pm$10.03 81.68$\pm$2.82 84.36$\pm$4.54 96.96$\pm$0.80 86.73$\pm$2.38 93.83$\pm$2.49 92.53$\pm$1.64
ELR 90.24$\pm$3.22 53.95$\pm$16.74 82.92$\pm$7.92 74.85$\pm$19.07 96.22$\pm$1.28 78.61$\pm$3.44 93.26$\pm$2.58 90.60$\pm$4.21
NCE-MAE 92.86$\pm$3.43 59.40$\pm$17.03 80.16$\pm$3.83 85.05$\pm$3.40 96.81$\pm$0.70 88.53$\pm$3.78 94.04$\pm$2.57 93.30$\pm$0.94
NCE-RCE 93.01$\pm$3.17 62.08$\pm$10.49 81.91$\pm$2.68 77.25$\pm$19.97 96.96$\pm$0.80 89.30$\pm$2.58 93.71$\pm$2.31 92.49$\pm$1.67
NFL-RCE 93.01$\pm$3.17 62.08$\pm$10.49 81.91$\pm$2.68 76.33$\pm$19.62 96.96$\pm$0.80 89.30$\pm$2.59 93.71$\pm$2.31 92.49$\pm$1.67
NCE-AGCE 93.06$\pm$3.24 62.16$\pm$10.52 81.96$\pm$2.68 77.30$\pm$20.00 96.94$\pm$0.78 89.49$\pm$2.38 93.74$\pm$2.33 93.87$\pm$2.32
NCE-AUL 92.97$\pm$3.13 62.00$\pm$10.42 81.87$\pm$2.67 84.27$\pm$4.59 96.89$\pm$0.72 89.13$\pm$2.64 93.80$\pm$2.52 92.36$\pm$1.56
NCE-AEL 93.04$\pm$3.32 65.16$\pm$6.39 80.33$\pm$4.00 83.95$\pm$4.50 96.52$\pm$0.42 87.70$\pm$3.87 93.97$\pm$1.49 93.78$\pm$0.77
ANL-CE 93.24$\pm$1.37 51.72$\pm$16.17 76.11$\pm$2.75 68.01$\pm$23.63 95.84$\pm$0.98 82.82$\pm$2.64 89.54$\pm$2.56 90.11$\pm$2.27
ANL-FL 92.72$\pm$0.99 49.91$\pm$14.62 73.46$\pm$6.29 73.51$\pm$18.85 95.96$\pm$1.13 82.46$\pm$2.41 86.96$\pm$2.25 88.90$\pm$1.74
ANN 97.42$\pm$0.33 86.36$\pm$3.06 91.34$\pm$1.50 92.63$\pm$2.37 98.24$\pm$0.23 91.57$\pm$1.84 93.22$\pm$1.39 95.60$\pm$0.85
WANN 97.29$\pm$0.38 89.81$\pm$2.73 92.51$\pm$2.40 93.77$\pm$2.29 98.18$\pm$0.18 93.72$\pm$1.77 95.41$\pm$0.69 96.84$\pm$0.40
Table 3: Accuracy ($\uparrow$) on limited noisy data settings. Both experiments are conducted on CIFAR-10, with $50$ noisy samples per class (left) and $100$ noisy samples per class (right). We highlight the significantly best method (paired $t$-test, $p<0.05$) or the top two methods if their difference is not statistically significant.
4.3 Limited noisy data
The inherent limitation of deep models lies in their reliance on a large amount of clean annotated data. In contrast, $k$-NN models demand less data, exhibiting potential robustness, particularly in the presence of label noise. To demonstrate this, two subsets are randomly subsampled from CIFAR-10, each containing only $50$ and $100$ samples per class. This experiment compares robust loss functions against ANN and WANN. As we do require a validation set for linear probing, we further extract $15\%$ of noisy samples to use as validation set, following previous experiments.
Artificial noise is introduced in CIFAR-10 using traditional methods: symmetric [35], asymmetric [39], and instance-dependent [50] noise. Symmetric noise involves the random switching of one label with any other label in the dataset. Asymmetric noise flips a given label with the adjacent one (i.e., label $i\mapsto i+1\mod C$, where $C$ is the number of classes). Finally, instance-dependent noise takes into account the relationships between the observations within the dataset. Following [58], we report the accuracy on 60% symmetric noise, 30% asymmetric noise, and 40% instance-dependent noise. Due to the stochasticity in subsampling and noise generation, we average the results across five seed runs.
Results   As shown in Table 3, across all experiments (i.e., two sample sizes and three noise injections each), WANN is clearly superior to both ANN and robust loss functions. The advantage of adaptive methods over robust loss functions is particularly evident under heavy noise conditions ($40\%$ inst. and $60\%$ asym.). Although this performance gap narrows slightly as noise rate decreases and data size increases (i.e., under less challenging conditions), the weighting scheme introduced in WANN continues to demonstrate clear benefits over ANN with minimal overhead. Moreover, a paired $t$-test ($p{<}0.05$) conducted across five individual runs confirms that WANN is the significantly best method under heavy noise conditions, underscoring its effectiveness in challenging scenarios. Figure 3 illustrates that under asymmetric noise, incorrect labels may cluster closely in the embedding space. This in fact poses challenges for the traditional $k$-NN but is mitigated by WANN. Notably, with 100 samples per class and $30\%$ asymmetric noise, while NCE-MAE approaches the performance of ANN, it does not surpass WANN. Additionally, while the difference between WANN and NCE-MAE is not statistically significant, WANN outperforms NCE-MAE by $1.37\%$ with an approximately four times lower standard deviation, suggesting potentially greater generalization capabilities.
Figure 3: t-SNE projection of a CIFAR-10 subset and its noisy (20% asymmetric) counterpart.
Figure 4: Average accuracy ($\uparrow$) and 95% confidence interval on stratified subsets of Animal-10N.
4.4 Limited real-world noisy data
We further assess classification performance on five stratified subsets of Animal-10N [41] consisting of $\{100,500,1000,2500,5000\}$ total samples. It is a widely adopted dataset for label noise benchmarks [54] and includes $50{,}000$ samples across $10$ classes, with approximately  ${\approx}\,8\%$ noisy labels. We compare the performance of ANN and WANN against three representative loss functions (cf. Figure 4). For linear probing, $15\%$ of the samples are retained as a validation set. No loss-hyperparameters were tuned for these datasets; instead, we used those found for CIFAR-10. For adaptive methods we adjust $(k_{\text{min}},k_{\text{max}})=(5,11)$ solely on the smallest sample size, as we expect  ${\approx}\,10$ samples per class. Accuracy is averaged across five runs.
Results   As shown in Figure 4, while linear probing approaches $k$-NN methods (i.e., ANN and WANN) with larger subsets, its effectiveness diminishes with limited data. Specifically, when reduced to about $10$ (noisy) samples per class, linear methods substantially underperform against WANN and ANN, exhibiting a much higher variance. This performance gap narrows only when the sample size is increased to about $2{,}500$ samples.
4.5 Generalization to medical data
To demonstrate the generalization capabilities and effectiveness of WANN, we provide another pool of experiments on two challenging medical datasets, which are semantically far from the pre-training data of the employed foundation model. Specifically, we used two datasets from the MedMNIST+ dataset collection [55], namely BreastMNIST and DermaMNIST, each with resolution $224{\times}224$. BreastMNIST is a binary dataset with only $546$ training samples. We limit our noise injection to symmetric noise at $20\%$, $30\%$, and $40\%$ noise ratios. In fact, on a binary dataset, symmetric and asymmetric noises are equal. DermaMNIST, on the other hand, is a multi-class dataset with $7$ classes and $7{,}007$ training samples. Thus, we apply the same previous noise settings (cf. Table 4). As before, we used the loss-hyperparameters defined for CIFAR-10. We report the average accuracy over five runs.
Results   Table 4 presents the results for the medical datasets. Notably, on BreastMNIST, WANN exhibits the highest accuracy with both clean and noisy training sets, highlighting its effectiveness in scenarios with limited data. Additionally, it significantly outperforms (paired $t$-test, $p<0.05$) ANN and robust loss functions under severe label noise (i.e., $40\%$ symmetric noise). While it might be expected that $k$-NN-based approaches would outperform linear methods on a binary problem with a small dataset, the weighting scheme in WANN demonstrates clear benefits on DermaMNIST as well. In fact, although linear methods outperform ANN, they do not surpass WANN, which consistently achieves a higher accuracy and lower standard deviation, demonstrating greater robustness.
BreastMNIST DermaMNIST
Pattern Clean Symmetric Symmetric Symmetric Clean Symmetric Asymmetric Instance
NR - 20% 30% 40% - 60% 30% 40%
[…]
Table 4: Accuracy ($\uparrow$) on BreastMNIST and DermaMNIST. We highlight the significantly best method (paired $t$-test, $p<0.05$) or the top two methods if their difference is not statistically significant.
4.6 Long-tailed noisy data
While $k$-NN is inherently robust, the selection of an appropriate value of $k$ is critical and not trivial, especially considering the unknown noise rate and potential class imbalances. In this experiment, we assess the performance of ANN and WANN, as compared to two fixed $k$-NN approaches. As discussed in the method section, both adaptive methods used $(k_{\text{min}},k_{\text{max}})=(11,51)$. They are compared with their respective fixed $k$-NN counterparts, namely $11$-NN and $51$-NN. For benchmarking on long-tailed problems, we use CIFAR-10LT and CIFAR-100LT, following prior studies [10, 14] with imbalance ratios of $1\%$ and $10\%$, respectively. It denotes the ratio between the least and most frequent class, with an exponentially decaying imbalance across all other classes. In CIFAR-10LT, the majority class has $5000$ samples, while the minority class consists of $50$ samples. Consequently, in CIFAR-100LT, we have $500$ and $50$ samples, respectively. As in previous experiments, noise is artificially injected. However, here we emphasize the generalizability across different noise patterns and severities. Thus, we inject symmetric noise ranging from $20\%$ to $60\%$, asymmetric noise from $20\%$ to $40\%$, and instance-dependent noise from $20\%$ to $40\%$, reporting the accuracy over five runs.
Results   The results presented in Figure 5 align with our expectations. Under limited noise conditions, a lower value of $k$ ($k=11$) generally achieves higher performance, while a higher value ($k=51$) shows greater robustness against more severe noise levels. In contrast, adaptive methods consistently approach the best performances across various noise rates and patterns. WANN’s weighting scheme consistently exhibits higher gains compared to ANN, especially under heavy asymmetric and instance-dependent noise. Additionally, considering all experiments and respective seed runs (due to the dataset variations from the stochastic noise injection), WANN emerges as the significantly best method, confirmed by the Wilcoxon signed-rank test ($p<0.05$).
((a)) Noisy CIFAR-10LT (1%)
((b)) Noisy CIFAR-100LT (10%)
Figure 5: Accuracy of fixed and adaptive $k$-NN based approaches on CIFAR-10LT and CIFAR-100LT, with $1\%$ and $10\%$ imbalance ratios. The imbalance ratio denotes the ratio between the least and the most frequent class, with an exponentially decaying imbalance between all other classes. Notably, WANN is close to the best performance across any noise pattern and severity, confirming its robustness and generalizability. Furthermore, WANN is the significantly best method across all experiments and datasets (Wilcoxon signed-rank test, $p<0.05$).
4.7 Dimensionality reduction
To showcase the effectiveness of the proposed Filtered LDA (FLDA) approach, we evaluate WANN’s classification accuracy on the linearly projected test set, utilizing the complete projected training set as WANN’s support set. Four different projections are employed for this analysis. First, as a baseline, we report classification performances without dimensionality reduction, labeled as “None”. Subsequently, we utilize Principal Component Analysis (PCA) with a fixed number of components, specifically $200$, explaining an average of $77.19\%$ of the total variance across each dataset. Additionally, we compare this with plain Linear Discriminant Analysis (LDA) fed with the entire noisy dataset. Finally, we present the effectiveness of our filtering approach before applying LDA, named as FLDA. To this extent, we recall that LDA projects the dataset along $C-1$ axes, where $C$ is the number of classes. Consistently with our previous experimental settings, the accuracy is reported on datasets with 60% symmetric noise, 30% asymmetric noise, and, 40% instance-dependent label noise. The experiments cover four different datasets: CIFAR-10, CIFAR-100 [25], MNIST [27], and Fashion-MNIST [51]. Note that all these datasets consist of 10 classes, except for CIFAR-100, which includes 100 classes.
Figure 6: Classification accuracy of WANN under three dimensionality reduction strategies, evaluated across four datasets and three different noise settings each. None reflects the performance using the original image embeddings without dimensionality reduction. PCA and LDA show the classification performances on the linearly projected feature spaces, utilizing the whole noisy training set. Finally, FLDA involves the proposed filtering approach. FLDA is the significantly best dimensionality reduction method across all experiments and datasets (Wilcoxon signed-rank test, $p<0.05$).
Results   The results are visually reported in Figure 6. As expected, PCA with $200$ components yields marginal improvement compared to the baseline, suggesting the need for a more nuanced linear projection. Surprisingly, LDA trained on the entire noisy label set already exhibits improved performance on some datasets relative to the baseline. This improvement could be attributed to the significantly reduced dimensionality, minimizing overall sparsity. Therefore, large neighborhoods may compensate for imperfect projections. However, FLDA is significantly better than all other techniques evaluated (Wilcoxon signed-rank test, $p{<}0.05$). The reduction in dimensionality not only enhances classification performance but also contributes to faster computational speed and potential storage efficiency. Notably, using DINOv2 ViT-L as feature extractor, we have an initial embedding size of $1024$. Thus, we significantly improved the classification performance with 100×100$\times$100 × and 10×10$\times$10 × smaller embeddings.
4.8 Explainability benefits
While deep learning methods are often viewed as black boxes with limited explainability, some traditional machine learning approaches inherently allow for a certain level of explainability. Notably, $k$-NN-based methods provide explainability through the representative nature of the neighborhood. Exploring the neighborhood for each prediction provides a deeper understanding of the factors contributing to correct or erroneous classifications. As illustrated in Figure 2, we present four test samples along with their top $3$ closest training examples from STL-10 [8], a subset of ImageNet [13]. In Figure 2(a), we highlight two test images with clearly incorrect reference labels, supported by clean training samples. In contrast, Figure 2(b) shows two ambiguous test labels, including multiple known objects within a single image, such as a bird with a ship in the background or a dog playing with a cat. The pitfalls of ImageNet labels were already illustrated by [4]. However, this provides a simple yet effective way to manually inspect potential wrong labels within our datasets. Indeed, the explainability of $k$-NN-based approaches becomes particularly valuable in the context of active label correction [36, 2], where human intervention is employed to re-annotate noisy samples. Beyond identifying noisy samples, these methods offer the potential to provide a reliable soft-label candidate derived from the majority of the neighborhood. Additionally, visualizing the neighbors can provide more evidence, especially for challenging or overlapping classes.
5 Discussion and conclusions
Limitations   The effectiveness of $k$-NN relies on a representative feature representation, posing challenges for domains not represented during pre-training. However, this limitation can be mitigated by using alternative feature extractors. This is now possible thanks to the growing availability of such large open-source models. Moreover, while the proposed FLDA significantly enhances classification performance, it requires a sufficient sample size to generate reliable projection vectors. Finally, to mitigate the computational demands of $k$-NN’s exhaustive search, established Approximate-NN methods [32, 17, 21] provide an effective solution, achieving comparable classification performance with substantially lower computational complexity. Indeed, we could expect a negligible influence on classification performance, as distance metrics inherently only approximate the nearest sample.
Side benefits and generalization   The efficiency of our classification method, further optimized with dimensionality reduction, enables its deployment with limited resources. Moreover, the reliable feature space of current feature extractors becomes particularly suitable for $k$-NN-based approaches, possibly generalizing to different domains. For instance, we observed outstanding classification and robustness performance in the medical domain, where interpretability, data scarcity, and data imbalance are still open challenges. Lastly, the compliance with recent regulations, such as the right to be forgotten [30], is facilitated by the proposed database approach. The flexibility to remove or add data points without extensive re-training or fine-tuning, aligns with privacy guidelines, providing a cost-effective solution.
Conclusion   We introduce WANN, a Weighted Adaptive Nearest Neighbor approach, relying on a reliability score quantifying the correctness of the training labels. Our experiments demonstrate its higher robustness with respect to label noise, particularly in scenarios with limited labeled data, surpassing traditional models relying on robust loss functions. Additionally, the incorporation of dimensionality reduction techniques enhances its performance under noisy labels, facilitated by the proposed weighted scoring.
Data Availability Statement
All experiments are conducted on publicly available datasets; see the references cited.
Acknowledgments
This study was funded through the Hightech Agenda Bayern (HTA) of the Free State of Bavaria, Germany.